### v1

# Hybrid Memetic Search for Electric Vehicle Routing with Time Windows, Simultaneous Pickup-Delivery, and Partial Recharges

Abstract
With growing environmental concerns, electric vehicles for logistics have gained significant attention within the computational intelligence community in recent years. This work addresses an emerging and significant extension of the electric vehicle routing problem (EVRP), namely EVRP with time windows, simultaneous pickup-delivery, and partial recharges (EVRP-TW-SPD), which has wide real-world applications. We propose a hybrid memetic algorithm (HMA) for solving EVRP-TW-SPD. HMA incorporates two novel components: a parallel-sequential station insertion procedure for handling partial recharges that can better avoid local optima compared to purely sequential insertion, and a cross-domain neighborhood search that explores solution spaces of both electric and non-electric problem domains simultaneously. These components can also be easily applied to various EVRP variants. To bridge the gap between existing benchmarks and real-world scenarios, we introduce a new, large-scale EVRP-TW-SPD benchmark set derived from real-world applications, containing instances with many more customers and charging stations than existing benchmark instances. Extensive experiments demonstrate the significant performance advantages of HMA over existing algorithms across a wide range of problem instances. Both the benchmark set and HMA will be open-sourced to facilitate further research in this area.
Index Terms:
Electric vehicle routing problem, memetic algorithm, combinatorial optimization, real-world application.
I Introduction
Over the past few decades, the vehicle routing problem (VRP) has drawn much attention in the computational intelligence community [1, 2, 3], due to its extensive real-world applications. As VRP research moves forward, a new focus has emerged: the electric vehicle routing problem (EVRP) [4]. This area has gained a lot of interest in recent years, as more and more electric vehicles (EVs) are being used in various sectors [5, 6, 7], such as public transit, home deliveries, postal services, distribution tasks, and emergency power supply. While EVs enable zero-emission logistics, operating a fleet of EVs presents additional challenges compared to operating conventional vehicles: i) the lower energy density of batteries compared to the fuel of combustion engine vehicles, necessitating EVs to perform longer detours to recharge batteries; ii) the limited number of public charging stations; and iii) the substantial time required for battery charging. To address these challenges, Erdogan and Miller-Hooks [8] introduced the first EVRP model that incorporated charging stations, as part of their broader study on the green VRP (G-VRP). Built on their work, subsequent studies have progressively incorporated more constraints into EVRP models to reflect real-world scenarios [9, 10].
This work focuses on an emerging and significant extension of EVRP, namely the electric vehicle routing problem with time windows, simultaneous pickup-delivery, and partial recharges (EVRP-TW-SPD), introduced by Akbay et al. [11]. Compared to the basic EVRP model, EVRP-TW-SPD considers more realistic problem characteristics and constraints, encompassing a wider range of real-world scenarios. Specifically, this model includes constraints such as limited driving range that requires en route recharging and strict customer service time windows [12]. Additionally, it incorporates two key characteristics: reverse logistics and partial recharges. In reverse logistics [13, 14], customers have both pickup and delivery demands, requiring EVs to complete both services in one trip. Partial recharges [15, 16] mean that EVs are not necessarily fully charged each time they visit a charging station, saving time to serve customers with tight time windows.
Despite the extensive real-world applications of EVRP-TW-SPD, current research in this area remains limited [11, 17, 18]. Specifically, existing approaches for solving EVRP-TW-SPD can be categorized into approximation methods [18] and metaheuristic methods [11, 17]. However, neither approach has shown fully satisfactory performance. Our experiments reveal that even for medium-scale instances with around 100 customers, there is significant room for improvement in solution quality obtained by existing methods. Furthermore, there is a notable gap between the scale of existing benchmark instances and real-world scenarios. The largest instance in the publicly available benchmark [11] contains only 100 customers and 21 charging stations. In contrast, with the rapid growth of urban areas, a real-world EVRP-TW-SPD instance might involve many more customers and stations [19, 20]. For example, application data from JD Logistics in Beijing  [^1] JD Logistics is one of the largest logistics companies in China., a major city in China, shows scenarios with up to 1000 customers and 100 charging stations. This disparity highlights the need for establishing larger-scale benchmarks to drive research that better addresses practical challenges.
This work aims to address the above research limitations. Specifically, we propose a novel hybrid memetic algorithm (HMA) for solving EVRP-TW-SPD. HMA combines individual-based large neighborhood search with population-based memetic search, leveraging their complementary strengths to quickly find high-quality solutions and further improve solution quality. Importantly, HMA introduces two key novelties. First, it incorporates a parallel-sequential station insertion procedure (PSSI) for handling partial recharges. Compared to the sequential station insertion procedure commonly used in the literature [15, 21, 22], PSSI can better avoid local optima and construct higher-quality routes with acceptable additional computational cost. Second, HMA employs a cross-domain neighborhood search (CDNS) as its local search module. CDNS simultaneously explores two different neighborhoods: those of the non-electric counterpart of EVRP-TW-SPD (i.e., VRP-TW-SPD) and EVRP-TW-SPD itself. This expanded search space not only leads to higher-quality solutions but also allows for easy reuse of efficient local search procedures from non-electric VRP variants, which are often more extensively studied. Finally, to address the gap in the scale of benchmark instances, we propose a new benchmark set based on real customer data from JD Logistics in the central area of Beijing.
The main contributions of this work are summarized below.
  1. 1.
    We propose a novel hybrid memetic algorithm (HMA) for solving EVRP-TW-SPD. HMA integrates two novel components: PSSI for handling partial recharges and CDNS as the local search module. We further discuss the potential of applying them to a wide range of EVRP variants beyond EVRP-TW-SPD.
  2. 2.
    We introduce a new, large-scale EVRP-TW-SPD benchmark set derived from real-world applications. This set, which will be made open-source, contains instances with up to 1000 customers and 100 charging stations, aiming to drive research in the field.
  3. 3.
    Extensive experiments validate the effectiveness of HMA, demonstrating its significant performance advantage over existing methods across a wide range of problem instances.
The rest of this paper is organized as follows. Section II presents a literature review. Section III formally defines the problem. Section IV first describes the novel components of HMA, followed by the hybrid framework. Section V compares the proposed algorithm against existing methods using the existing benchmark and introduces the new benchmark. Finally, Section VI concludes the paper.
II Related Work
In this section, we first review literature related to two key aspects of EVRP-TW-SPD: reverse logistics and partial recharges, followed by existing methods for the problem.
Reverse logistics, which involves the bi-directional flow of goods in pickup and delivery, is crucial in modern transportation due to its importance in satisfying distinct customer demands and lowering energy consumption. According to [23], the VRPs involving reverse logistics have often been referred to as pickup and delivery problems (PDPs). Among PDPs, the VRP with simultaneous pickup-delivery (VRPSPD) is the most studied variant [24, 25], where vehicles deliver goods from a central depot to customers while also picking up items from customers to return to the depot. VRPSPD was first studied in [26] for a book distribution system with 22 customers and two vehicles, demonstrating that combining pickup and delivery operations leads to significant savings in time and distance compared to separate operations. Since then, numerous studies have been conducted on VRPSPD and its variants. In EVRP-TW-SPD, to handle simultaneous pickup-delivery, existing techniques from VRPSPD research [27, 28, 13] can be adapted. Specifically, in this work we integrate the residual capacity and radical surcharge (RCRS) [13] construction heuristic into the HMA algorithm.
Compared to full recharges, partial recharges are a more realistic and practical charging scheme in the real world. It allows for flexible charging time, which can help EVs meet tight time windows and potentially serve more customers. To handle partial recharges, two key aspects need to be considered: station selection and charge amount calculation. First, station selection refers to choosing which charging station EVs should use when they need to recharge. Most existing approaches for station selection rely on heuristic insertion strategies. Some approaches, like those in [29, 30], focus on minimizing the distance from the current customer to the nearest charging station. Others, such as [11, 17, 15, 31], aim to minimize the increase in total distance caused by inserting a charging station. Specifically, Keskin et al. [15] proposed three station insertion heuristics: greedy station insertion, greedy station insertion with comparison, and best station insertion. In contrast, Wang et al. [20] took a different approach by minimizing a generalized cost function that includes the objective function, time window, and distance violations. Second, charge amount calculation involves determining how much to recharge at each charging station. Several approaches have been proposed. Felipe et al. [32] suggested charging only the minimum amount needed to reach the next charging station or depot. Keskin et al. [15] ensured sufficient charge to reach the next station or depot while adjusting for time window constraints. They also found that partial recharges perform significantly better than full recharges, even with pre-determined charging amounts. Desaulniers et al. [16] and Wang et al. [20] considered both the minimum required charge and available slack time when determining charge amount. Similar to [16, 20], our approach for determining charge amount also considers both the minimal required charge and available slack time, ensuring sufficient battery state while meeting all time constraints.
Currently, research on EVRP-TW-SPD is still rather limited. Akbay et al. [11] introduced the problem and formulated it as a mixed integer linear programming (MILP) model. They proposed a metaheuristic approach called Adapt-CMSA-STD and compared it with CPLEX and other heuristics on instances with up to 100 customers and 21 charging stations. Their follow-up work [17] presented Adapt-CMSA-SETCOV, which used vehicle-route representations and a set-covering-based MILP model and showed performance improvement over Adapt-CMSA-STD. Feng et al. [18] developed an alternating direction multiplier method for EVRP-TW-SPD with full recharges, testing it on small-scale instances with up to 15 customers and 6 charging stations.
III Notations and Problem Definition
Formally, the EVRP-TW-SPD is defined on a complete directed graph $\bm{G}=(\bm{V}^{\prime},\bm{E})$, where $\bm{V}^{\prime}=\{0\}\cup\bm{V}\cup\bm{F}=\{0,1,2,\ldots,M+P\}$ is the node set. Here, 00 represents the depot, $\bm{V}=\{1,2,\ldots,M\}$ is a set of $M$ customers, and $\bm{F}=\{M+1,\ldots,M+P\}$ is a set of $P$ charging stations allowing multiple visits. The arc set $\bm{E}=\{(i,j)\mid i,j\in\bm{V}^{\prime},i\neq j\}$ is defined between each pair of nodes. Each node $i\in\bm{V}^{\prime}$ has five attributes: delivery demand $u_{i}$, pickup demand $v_{i}$, time window $[e_{i},l_{i}]$, and service time $s_{i}$. For a customer $i$, $u_{i}$ is the amount of goods delivered from the depot, and $v_{i}$ is the amount of goods picked up to be returned to the depot. The time window $[e_{i},l_{i}]$ defines the earliest and latest time for customer service, with arrivals before $e_{i}$ resulting in waiting and after $l_{i}$ being infeasible. $s_{i}$ is the unloading/loading time at customer $i$. For the depot 00, $e_{0}$ and $l_{0}$ are the earliest departure and latest return time for EVs, respectively, with $u_{0}=v_{0}=s_{0}=0$. For a charging station $i\in\bm{F}$, $u_{i}=v_{i}=s_{i}=0$, and $e_{i}=e_{0}$, $l_{i}=l_{0}$. Each arc $(i,j)\in\bm{E}$ is associated with a travel distance $d_{ij}$ and a travel time $t_{ij}$.
A homogeneous fleet of EVs, all with identical loading capacity $C$, battery capacity $Q$, and dispatching cost $\mu_{1}$, is initially located at the depot. The EVs depart from the depot, serve customers, and finally return to the depot. When an EV visits a charging station, its battery is recharged at a constant rate of $g>0$, i.e., it takes time $g$ to charge one unit of battery energy. Conversely, the constant $h$ represents the EV’s battery energy consumption rate per unit of travel distance for the EV.
A solution $\bm{S}$ to EVRP-TW-SPD is represented by a set of vehicle routes, $\bm{S}=\{\bm{R}_{1},\bm{R}_{2},\ldots,\bm{R}_{K}\}$, where $K$ is the number of vehicles used. Each route $\bm{R}_{i}$ consists of a sequence of nodes that the EV visits, $\bm{R}_{i}=(n_{i,1},n_{i,2},\ldots,n_{i,L_{i}})$, where $n_{i,j}$ is the $j$-th node visited in $\bm{R}_{i}$, and $L_{i}$ is the length of $\bm{R}_{i}$. For brevity, we temporarily omit the subscript $i$ in $\bm{R}_{i}$ in the following, i.e., $\bm{R}=(n_{1},n_{2},\ldots,n_{L})$. The total travel distance of $\bm{R}$, denoted as $TD(\bm{R})$, is:
$TD(\bm{R})$ $=\sum_{j=1}^{L-1}d_{n_{j}n_{j+1}}.$ (1)
The battery energy state of the EV on arrival at and departure from $n_{j}$, denoted as $y_{n_{j}}$ and $Y_{n_{j}}$, respectively, can be computed recursively as follows:
$y_{n_{j}}$ $=Y_{n_{j-1}}-h\cdot d_{n_{j-1}n_{j}},\ j>1,$ (2)
$Y_{n_{j}}$ =$=\begin{cases}Q,&\ j=1,\\ y_{n_{j}},&\ j>1,\ n_{j}\notin\bm{F},\\ y_{n_{j}}+q_{n_{j}},&\ j>1,\ n_{j}\in\bm{F},\end{cases}$
where $q_{n_{j}}$ is the charge amount at charging station $n_{j}\in\bm{F}$. An EV departs from the depot with a full battery. The arrival and departure time at $n_{j}$, denoted as $a_{n_{j}}$ and $b_{n_{j}}$, respectively, can be computed recursively as follows:
$a_{n_{j}}$ $=b_{n_{j-1}}+t_{n_{j-1}n_{j}},\ j>1,$ (3)
$b_{n_{j}}$ =$=\begin{cases}e_{0},&\ j=1,\\ \max\{a_{n_{j}},e_{n_{j}}\}+s_{n_{j}},&\ j>1,\ n_{j}\notin\bm{F},\\ a_{n_{j}}+g\cdot q_{n_{j}},&\ j>1,\ n_{j}\in\bm{F}.\\ \end{cases}$
The EV load on arrival at $n_{j}$, denoted as $load_{n_{j}}$, is:
$load_{n_{j}}$ =$=\begin{cases}\sum_{j=1}^{L}u_{n_{j}},&\ j=1,\\ load_{n_{j-1}}-u_{n_{j-1}}+v_{n_{j-1}},&\ j>1.\end{cases}$ (4)
The total cost of $\bm{S}$, denoted as $TC(\bm{S})$, consists of two parts: the dispatching cost, $\mu_{1}\cdot K$, and the transportation cost, which is the total travel distance of $\bm{S}$ multiplied by the cost per unit of travel distance $\mu_{2}$. The objective is to find a solution with the minimum $TC$, as presented in Eq. (5a):
$\min_{\bm{S}}TC(\bm{S})=\mu_{1}\cdot K+\mu_{2}\cdot\sum_{i=1}^{K}TD(\bm{R}_{i})$ (5a)
[…]
$\quad b_{n_{i,1}}\geq e_{0}\ \text{and}\ a_{n_{i,L_{i}}}\leq l_{0},\ 1\leq i\leq K,$ (5f)
[…]
Constraint (5b) ensures that each route starts from and returns to the depot. Constraint (5c) ensures that each customer is served exactly once (note $\mathbb{I}[\cdot]$ is the indicator function). Constraint (5d) ensures that the delivery and pickup demands of customers are met simultaneously while EVs do not exceed their loading capacity $C$. Constraint (5e) requires EVs to visit each node within its corresponding time window. Constraint (5f) ensures that EVs depart only after the start of the depot’s time window and return before the end of its time window. Finally, constraint (5g) requires EVs to maintain a non-negative battery state at all nodes; when serving customers or visiting charging stations, the battery level will not decrease, and it will not exceed battery capacity $Q$ when leaving any node.
In this work, unless otherwise specified, we evaluate solutions based on their cost (as their fitness) defined in Eq. (5a), where a lower cost indicates a better solution.
IV Hybrid Memetic Algorithm for EVRP-TW-SPD
In this section, we first introduce the two novel components integrated into HMA: PSSI for handling partial recharges and CDNS as the local search module. Then, we describe the overall framework of HMA.
IV-A Parallel-Sequential Station Insertion (PSSI)
To transform electricity-infeasible routes into high-quality feasible routes, we propose a parallel-sequential station insertion (PSSI) procedure. Compared to existing sequential insertion procedures [15, 21, 22] that typically lead to suboptimal routes, PSSI can better avoid local optima and generate higher-quality routes with acceptable additional computational cost. As shown in Algorithm 1, PSSI leverages two distinct strategies: parallel station insertion (PSI) and sequential station insertion (SSI). PSI (lines 1-1) employs a genetic algorithm to search for optimal station insertions between all consecutive customer pairs simultaneously, determining both the necessity of insertions and the inserted stations. SSI (lines 1-1), on the other hand, inserts charging stations sequentially along the route. Finally, PSSI selects the better route between PSI and SSI as the final outcome (line 1). By combining PSI and SSI, PSSI enables a more comprehensive exploration of the route space, potentially yielding higher-quality routes compared to purely sequential procedures.
As aforementioned in Section II, handling partial recharges involves two key aspects: station selection and charge amount calculation. In terms of station selection, PSI is designed for scenarios where at most one charging station is needed between any two consecutive nodes, while SSI allows multiple charging stations to be inserted between two consecutive nodes. Regarding charge amount calculation, both PSI and SSI use the same method. Below, we detail PSI and SSI based on these two aspects.
Input: an electricity-infeasible route: $\bm{R}$
Output: the best-found feasible route: $\bm{R}^{*}$
1
/* Parallel Station Insertion (PSI) */
2 $\bar{\bm{R}}\leftarrow$ remove all stations in $\bm{R}$;
3 $\{(\bm{x}_{1},\bm{R}_{1}),\ldots,(\bm{x}_{\alpha L},\bm{R}_{\alpha L})\}\leftarrow$ Initialization($\bar{\bm{R}},\alpha$);
4 while max generation count $B$ not reached do
5       for $i\leftarrow 1$ to $\alpha L$ do
6             $\bm{x}_{\text{p}_{1}},\bm{x}_{\text{p}_{2}}\leftarrow\text{Random\_Select}(\bm{x}_{1},\bm{x}_{2},\ldots,\bm{x}_{\alpha L})$;
7             $\bm{x}_{\text{child}}\leftarrow\bm{x}_{\text{p}_{1}}\oplus\bm{x}_{\text{p}_{2}}$;
8             $\bm{x}_{\text{child}}\leftarrow\text{Mutation}(\bm{x}_{\text{child}})$;
9             $\bm{R}_{\text{child}}\leftarrow$ insert stations into $\bar{\bm{R}}$ according to $\bm{x}_{\text{child}}$;
10             if $\bm{R}_{\mathrm{child}}$ is better than $\bm{R}_{i}$ then $(\bm{x}_{i},\bm{R}_{i})\leftarrow(\bm{x}_{\text{child}},\bm{R}_{\text{child}})$;
11
12       end for
13
14 end while
15 $\bm{R}_{\text{PSI}}\leftarrow\text{Select\_Best}(\bm{R}_{1},\ldots,\bm{R}_{\alpha L})$;
16
/* Sequential Station Insertion (SSI) */
17 $\bm{R}^{\prime}\leftarrow\bm{R}$;
18 while $\bm{R}^{\prime}$ is electricity-infeasible do
19       $n_{\text{right}}\leftarrow$ the first node with negative battery state in $\bm{R}^{\prime}$;
20       $n_{\text{left}}\leftarrow$ the last station or depot visted before $n_{\text{right}}$ in $\bm{R}^{\prime}$;
21       $\bm{R}^{\prime}\leftarrow$ perform the best-possible station insertion in the path from $n_{\text{left}}$ to $n_{\text{right}}$;
22
23 end while
24 $\bm{R}_{\text{SSI}}\leftarrow$ improve consecutive stations in $\bm{R}^{\prime}$;
25
[…]
27
return $\bm{R}^{*}$
Algorithm 1 PSSI
IV-A1 Station Selection
For a given problem instance $\bm{G}=(\bm{V}^{\prime},\bm{E})$, we preprocess the charging stations to rank them for insertion between each pair of nodes. The ranking metric is the extra cost induced by insertion, formally defined as $d_{ik}+d_{kj}-d_{ij}$, where $k$ is a potential charging station to be inserted between nodes $i$ and $j$. To balance efficiency and solution quality, only the top $sr\cdot|\bm{F}|$ ranked charging stations are considered for insertion in both PSI and SSI, where $sr\in(0,1]$ is the selection range parameter.
PSI utilizes a genetic algorithm that consists of two main stages: initialization (lines 1-1) and evolution (lines 1-1). In the initialization stage, all charging stations are first removed from the input electricity-infeasible route $\bm{R}$, resulting in a fixed route $\bar{\bm{R}}=(c_{0},c_{1},c_{2},\ldots,c_{L-1},c_{L})$, where $c_{0}=c_{L}=0$ represent the depot, and $c_{1},c_{2},\ldots,c_{L-1}$ are the $L-1$ customers. We define a binary decision variable $X_{j,j+1}$ to indicate whether a charging station is inserted between $c_{j}$ and $c_{j+1}$ (1 for insertion, 0 for no insertion). Thus, the station insertion information is represented as a binary sequence $\bm{x}=(X_{0,1},X_{1,2},\ldots,X_{L-1,L})$. An initial population of $\alpha L$ binary sequences (i.e., the population size is proportional to $L$, with $\alpha$ representing the proportionality parameter) is first generated, where each $X_{j,j+1}$ is randomly set to 1 or 0 with equal probability (line 1). For each sequence, the binary encoding determines whether to insert a station between two consecutive nodes, and we always attempt to insert the highest-ranked station when insertion is indicated. If this insertion violates constraints, we try the next highest-ranked station until a feasible insertion is found or all stations within the selection range have been attempted. This process results in a set of feasible routes $\bm{R}_{1},\bm{R}_{2},\ldots,\bm{R}_{\alpha L}$.
In the evolution stage, the population evolves over $B$ generations. In each generation, two parent sequences are first randomly selected. A child sequence $\bm{x}_{\text{child}}$ is generated using bitwise XOR crossover (denoted by $\oplus$), inheriting unique insertion positions from each parent. Then, mutation is applied to $\bm{x}_{\text{child}}$: each bit has a low chance of flipping (i.e., 2%), followed by a high chance of flipping from 1 to 0 (i.e., 20%) to encourage fewer station insertions. The corresponding route $\bm{R}_{\text{child}}$ is evaluated. Finally, a sequence in the population is replaced with $\bm{x}_{\text{child}}$ if $\bm{R}_{\text{child}}$ is with higher quality (line 1). The best route $\bm{R}_{\text{PSI}}$ among the final population is selected as the outcome of PSI (line 1).
Unlike PSI, SSI progressively inserts stations along the route (lines 1-1). Specifically, for the first node $n_{\text{right}}$ in the route with a negative battery level, it first finds the last charging station or depot visited before this node, denoted as $n_{\text{left}}$. Then, it evaluates all potential station insertions in the path from $n_{\text{left}}$ to $n_{\text{right}}$. Finally, SSI selects the best feasible insertion to perform, i.e., the one that increases the route distance the least (lines 1-1). However, this greedy procedure can be short-sighted, particularly when multiple stations are inserted between two consecutive nodes. To address this limitation, the SSI procedure includes an additional refinement step (line 1). That is, for consecutive stations in the route, we attempt to replace them with stations based on the preprocessed rankings, starting from the highest-ranked to lower-ranked stations. If an improvement is found, the replacement is made. The final outcome of the SSI procedure is $\bm{R}_{\text{SSI}}$.
IV-A2 Charge Amount Calculation
We propose a new procedure tailored to EVRP-TW-SPD for calculating the partial recharge amount, building upon the work of [16, 20]. Our approach considers both the minimal charge amount required and the available slack time, ensuring that the battery state is sufficient to reach the next charging station or depot while satisfying all time window constraints.
We define a path $\bm{P}_{i}=(f_{i},c_{i,1},c_{i,2},\ldots,c_{i,m_{i}},f_{i+1})$ in route $\bm{R}$, where $f_{i}$ and $f_{i+1}$ are two successive charging stations (or the depot) in the route, separated by $m_{i}$ customers $c_{i,j}\ (j=1,...,m_{i})$. Our approach aims to minimize waiting time by utilizing any available slack time for additional charging. The charge amount at $f_{i}$ consists of two parts: the minimal charge $q_{f_{i},0}$ and the additional charge $q_{f_{i},1}$. Specifically, $q_{f_{i},0}$ is calculated as:
$q_{f_{i},0}$ $=\max\Bigg{\{}0,h\cdot TD(\bm{P}_{i})-y_{f_{i}}\Bigg{\}},$ (6)
where $h$ is the energy consumption rate and $TD(\bm{P}_{i})$ is the total distance of path $\bm{P}_{i}$.
To compute the additional charge $q_{f_{i},1}$, we define the following variables. Let $a^{\prime}_{c_{i,j}}$, $b^{\prime}_{c_{i,j}}$ be the updated arrival and departure time at $c_{i,j}$ (distinguished from $a_{c_{i,j}}$, $b_{c_{i,j}}$ in Section III), $\delta_{i,j}$ be the slack time that can be used to charge due to waiting before starting service at $c_{i,j}$, $\tau_{i,j}$ be the potential available time that can be used to charge when leaving $c_{i,j}$ while satisfying all previous time window constraints. Without loss of generality, $f_{i}$ is regarded as $c_{i,0}$ for notation convenience. Hence, $q_{f_{i},1}$ can be computed as:
$q_{f_{i},1}=\frac{1}{g}\sum_{j=1}^{m_{i}}\delta_{i,j},$ (7)
where $g$ is the charging rate and $\delta_{i,j}$ can be computed recursively as follows:
$a^{\prime}_{c_{i,j}}$ $=b^{\prime}_{c_{i,j-1}}+t_{c_{i,j-1}c_{i,j}},\ j>0,$ (8)
$b^{\prime}_{c_{i,j}}$ =$=\begin{cases}a_{f_{i}}+g\cdot q_{f_{i},0},&\ j=0,\\ \max\{e_{c_{i,j}},a^{\prime}_{c_{i,j}}+\delta_{i,j}\}+s_{c_{i,j}},&\ j>0,\\ \end{cases}$
$\delta_{i,j}$ $=\min\{\tau_{i,j-1},\max\{0,e_{c_{i,j}}-a^{\prime}_{c_{i,j}}\}\},\ j>0,$
$\tau_{i,j}$ =$=\begin{cases}\infty,&\ j=0,\\ \min\{\tau_{i,j-1},l_{c_{i,j}}-a^{\prime}_{c_{i,j}}\}-\delta_{i,j},&\ j>0.\\ \end{cases}$
When $\tau_{i,j}$ is 00, the recursive equations can terminate prematurely. Upon reaching $c_{i,m_{i}}$ of $\bm{P}_{i}$, if no slack time is available, even if $\tau_{i,m_{i}}>0$, no more charge time is needed. Thus, the additional charge $q_{f_{i},1}$ does not affect the arrival time at the next station $f_{i+1}$.
Finally, the total charge amount $q_{f_{i}}$ that does not exceed $Q-y_{f_{i}}$ when visiting $f_{i}$ is computed as:
$q_{f_{i}}$ $=\min\{q_{f_{i},0}+q_{f_{i},1},Q-y_{f_{i}}\}.$ (9)
IV-A3 Discussion
Compared to existing purely sequential insertion procedures [15, 21, 22], PSSI can typically obtain higher-quality feasible routes with reasonable additional computational cost, as demonstrated in the experiments (see Section V-D). The applicability of PSSI extends beyond EVRP-TW-SPD to various EVRP variants sharing certain features: i) partial recharges, ii) customers with time windows, and iii) a homogeneous network of charging stations. For instance, PSSI could be applied to scenarios involving multiple depots, heterogeneous vehicle fleets, or even dynamic routing where customer demands or traffic conditions change in real-time.
IV-B Cross-Domain Neighborhood Search (CDNS)
We propose cross-domain neighborhood search (CDNS), a novel local search procedure that explores two distinct solution spaces simultaneously: the solution space of EVRP-TW-SPD and the solution space of its non-electric counterpart VRP-TW-SPD [19], where some solutions may be infeasible for EVRP-TW-SPD. This cross-domain exploration offers two key advantages. First, by expanding beyond the feasible solution space of EVRP-TW-SPD, CDNS potentially discovers higher-quality solutions that might be overlooked by conventional single-domain local search. Second, exploring the solution space of VRP-TW-SPD enables the efficient reuse of well-established local search procedures for VRP-TW-SPD. We will further discuss the broader implications of this cross-domain approach beyond EVRP-TW-SPD at the end of Section IV-B.
The core idea of CDNS is to exploit the relationship between feasible solutions of EVRP-TW-SPD and VRP-TW-SPD [19], as stated below.
Proposition 1.
Given an EVRP-TW-SPD instance, removing all charging stations from a feasible EVRP-TW-SPD solution always results in a feasible VRP-TW-SPD solution.
The proof can be found in Appendix A. Conversely, to transform a feasible VRP-TW-SPD solution to a feasible EVRP-TW-SPD solution, we can use the aforementioned PSSI procedure.
Input: a feasible solution: $\bm{S}$
Output: a local optimum: $\bm{S}$
/* Aggressive Local Search (ALS) */
1 $\bm{S}^{\prime}\leftarrow\bm{S}$;
2 $\bar{\bm{S}}\leftarrow$ remove all stations in $\bm{S}^{\prime}$;
3 repeat
4       $\bar{\bm{S}}\leftarrow$ apply an existing local search procedure for VRP-TW-SPD to find the best-improving solution in the neighborhood of $\bar{\bm{S}}$;
5       for each route $\bar{\bm{R}}\in\bar{\bm{S}}$ do
6             $\bm{R}\leftarrow\text{PSSI}(\bar{\bm{R}})$;
7             replace the corresponding route in $\bm{S}^{\prime}$ with $\bm{R}$;
8
9       end for
10       if $\bm{S}^{\prime}$ is better than $\bm{S}$ then $\bm{S}\leftarrow\bm{S}^{\prime}$;
11
12until $\bar{\bm{S}}$ is not improved or $\bm{S}^{\prime}$ is electricity-infeasible;
13
/* Conservative Local Search (CLS) */
14 repeat
15       $\bm{S}\leftarrow$ find the best-improving solution in the EVRP-TW-SPD neighborhood of $\bm{S}$;
16
17until $\bm{S}$ is not improved;
return $\bm{S}$
Algorithm 2 CDNS
Algorithm 2 outlines the framework of CDNS. It consists of two sub-procedures. The first, called aggressive local search (ALS), explores the solution space of VRP-TW-SPD. As its name suggests, it temporarily ignores the electricity constraint (5g) and repeatedly applies existing local search procedures for VRP-TW-SPD to find the solution with maximum improvement (i.e., the best-improving solution) in the neighborhood of the current solution (lines 2-2). When ALS can no longer improve the solution, CDNS switches to conservative local search (CLS) which searches for improvements within the solution space of EVRP-TW-SPD (lines 2-2). Below we detail these two sub-procedures.
IV-B1 Aggressive Local Search (ALS)
When a feasible solution $\bm{S}$ to EVRP-TW-SPD enters ALS (lines 2-2), the first step is to remove all charging stations in $\bm{S}$, resulting in a feasible solution $\bar{\bm{S}}$ to VRP-TW-SPD (line 2). Then, an existing local search procedure for VRP-TW-SPD is repeatedly applied to find the best-improving solution in the neighborhood of $\bar{\bm{S}}$ (line 2). Here, we employ the local search procedure of MATE [19], one of the state-of-the-art algorithms for solving VRP-TW-SPD. This procedure involves multiple move operators and constant-time move evaluation (see the paper of MATE for details). Then, for each found best-improving solution, PSSI is first applied to transform it into a feasible EVRP-TW-SPD solution (lines 2-2). If this transformation is not possible, then ALS would terminate (i.e., the second termination condition in line 2). If successful, the resulting feasible solution would replace the current solution $\bm{S}$ when it has better quality (line 2). ALS would terminate when no further improvement is possible (i.e., the first termination condition in line 2).
IV-B2 Conservative Local Search (CLS)
Unlike ALS, CLS performs local search directly in the solution space of EVRP-TW-SPD. It repeatedly identifies the best-improving solution in the current solution’s neighborhood until no further improvement is possible (lines 2-2). The neighborhood is defined by the following five move operators.
  • 2-opt: it inverts a subsequence of two consecutive customers in a route.
  • 2-opt*: it removes two arcs from two different routes, splitting each into two parts, then reconnects the first part of one route with the second part of the other, and vice versa.
  • or-opt: it removes one or two consecutive customers from a route and reinserts them elsewhere in the same or a different route.
  • swap: it exchanges two non-overlapping subsequences of one or two consecutive customers, either within the same route or across different routes.
  • relocate: it moves a customer from its current position to a new position, either within the same route or to a different route.
IV-B3 Discussion
By exploring two distinct solution spaces, CNDS typically finds higher-quality solutions than using either solution space alone, as demonstrated in our experiments (see Section V-D). The idea of CDNS can be readily generalized to other EVRP variants by simultaneously exploring both electric and non-electric neighborhoods. This dual-domain approach is particularly valuable since non-electric variants are typically more extensively studied in the literature, providing a rich pool of well-established local search procedures that can be adapted for solving various EVRPs efficiently.
IV-C Overall Framework of HMA
Figure 1: The flow chart of HMA.
We now introduce the framework of HMA, which integrates an individual-based large neighborhood search (LNS) and a population-based memetic search (MS). The flow chart of HMA is illustrated in Figure 1. Overall, HMA employs LNS to efficiently obtain a high-quality solution, and then uses MS to further improve it.
As presented in Algorithm 3, HMA first enters the individual-based LNS. It begins with the initialization of an individual solution $\bm{S}$, which is constructed using the RCRS heuristic [13]. Specifically, RCRS is an extension of the cheapest-insertion heuristic and constructs routes by iteratively inserting an unassigned customer at the position with the minimal value according to the RCRS criterion until no feasible insertions are possible due to capacity or time window constraints (please refer to [13] for details of RCRS). The RCRS criterion combines the total travel distance, residual capacity, and radial surcharge with weighting parameters $\lambda,\gamma\in[0,1]$. Here, we simply set $\lambda=\gamma=0.5$. In this initialization, electricity-feasible customers are first considered for insertion, followed by those that are only electricity-infeasible but can be handled later through best station insertion. When no feasible insertions remain, a new route is initiated. This process continues until all customers are assigned (line 3).
Input: an EVRP-TW-SPD instance: $\bm{G}=(\bm{V}^{\prime},\bm{E})$
Output: the best-found solution: $\bm{S}$
1
/* Large Neighborhood Search (LNS) */
2 $\{\bm{S}\}\leftarrow\text{Initialization}()$;
3 while true do
4       $\bm{S}^{\prime}\leftarrow\bm{S}$;
5       remove $\rho$ customers from $\bm{S}^{\prime}$ and reinsert them into $\bm{S}^{\prime}$ via regret-based insertion;
6       $\bm{S}^{\prime}\leftarrow\text{CDNS}(\bm{S}^{\prime})$;
7       if $\bm{S}^{\prime}$ is better than $\bm{S}$ then $\bm{S}\leftarrow\bm{S}^{\prime}$;
8
9      if $\bm{S}$ is not improved in the last $G$ iterations then
             /* Memetic Search (MS) */
10             $\bm{S}^{\prime}\leftarrow\bm{S}$;
11             $\{\bm{S}_{1},\ldots,\bm{S}_{N}\}\leftarrow\text{Population\_Initialization}(\bm{S}^{\prime})$;
12             repeat
13                   $\pi(\cdot)\leftarrow\text{a random permutation of }1,\ldots,N$;
14                   for $i\leftarrow 1$ to $N$ do
15                         $\bm{p}_{1}\leftarrow\bm{S}_{\pi(i)},\ \bm{p}_{2}\leftarrow\bm{S}_{\pi(i+1)}$;
16                         $\bm{S}_{\text{child}}\leftarrow\text{Crossover}(\bm{p}_{1},\bm{p}_{2})$;
17                         $\bm{S}_{\text{child}}\leftarrow$ CDNS$(\bm{S}_{\text{child}})$;
18                         $\bm{S}_{\pi(i)}\leftarrow\text{Select\_Best}(\bm{S}_{\text{child}},\bm{p}_{1})$;
19
20                   end for
21                   $\bm{S}^{\prime}\leftarrow\text{Select\_Best}(\bm{S}^{\prime},\bm{S}_{1},\ldots,\bm{S}_{N})$;
22
23            until $\bm{S}^{\prime}$ is not improved in the last $G$ generations;
24             if $\bm{S}^{\prime}$ is better than $\bm{S}$ then $\bm{S}\leftarrow\bm{S}^{\prime}$;
25             else break;
26
27       end if
28
29 end while
return $\bm{S}$
Algorithm 3 HMA
After initialization, HMA applies the destroy-repair operator as described in [19] (a large-step perturbation) on the individual solution $\bm{S}$ (line 3). This destroy-repair operator involves removing $\rho$ correlated customers from the solution, where the random number $\rho\in[\omega_{1}|\bm{V}|,\omega_{2}|\bm{V}|]$, and $\omega_{1}$, $\omega_{2}$ are parameters with $10<\omega_{1}\leq\omega_{2}<$ 1. The operator then reinserts these customers using a regret-based heuristic that evaluates the expected cost of inserting a node in a future iteration rather than in the current one to assess the quality of potential insertions. The process is followed by CDNS, ensuring that a new local optimum is reached to potentially improve $\bm{S}$ (lines 3-3). If $\bm{S}$ does not show improvement over a specified number of consecutive iterations, denoted as $G$, HMA initiates the population-based MS (lines 3-3).
MS starts with a population of $N$ individuals $\bm{S}_{1},...,\bm{S}_{N}$, generated based on the current elite solution $\bm{S}$ (line 3). Specifically, $\bm{S}_{1}$ is set as a copy of $\bm{S}$ because the elite solution should be preserved. As for the remaining individuals, the individual $\bm{S}_{i}$ with odd index $i$ comes from the destroy-repair operator performed on $\bm{S}$ with $\omega_{1}=\omega_{2}=i/N$; for the individual $\bm{S}_{i}$ with even index $i$, it is generated by RCRS with $\lambda=\frac{1}{\sqrt{N}-1}\cdot(p-1),\gamma=\frac{1}{\sqrt{N}-1}\cdot(q-1)$, where $p=\left\lceil i/\sqrt{N}\right\rceil$ and $q=i-\sqrt{N}\cdot(p-1)$. The above procedure would result in a diverse population of individuals while preserving the high-quality information of the elite solution obtained from the previous LNS.
After initialization, MS enters an evolutionary process (lines 3-3). In each generation (lines 3-3), each individual in the population is selected once as parent $\bm{p}_{1}$ and once as parent $\bm{p}_{2}$, in a random order (lines 3 and 3). For each pair of parent solutions, MS generates an offspring using the route-assembly-regret-insertion (RARI) crossover [19] (line 3), and then improves it through CDNS (line 3). After that, the better one among the offspring and $\bm{p}_{1}$ replaces the individual selected as $\bm{p}_{1}$ (line 3). After MS, a new candidate solution $\bm{S}^{\prime}$ is obtained with the aim of improving $\bm{S}$ (lines 3-3). If the quality of $\bm{S}$ is still not improved by MS, then HMA would output $\bm{S}$ and terminate (line 3); otherwise, HMA proceeds to the next iteration of LNS (line 3).
When presented with a large-scale problem instance (with more than 200 customers) to solve, we make two adjustments to HMA to improve its efficiency. First, only ALS is used in CDNS. Second, problem decomposition is employed in MS. Specifically, the decomposition breaks down the main problem $\bm{G}=(\bm{V}^{\prime},\bm{E})$ into $k$ smaller subproblems $\bm{G}_{1},\ldots,\bm{G}_{k}$ when entering MS (line 3). A subproblem $\bm{G}_{i}=(\bm{V}_{i}^{\prime},\bm{E}_{i})$ with $\bm{V}_{i}^{\prime}=\{0\}\cup\bm{V}_{i}\cup\bm{F}$, $\bm{E}_{i}=\{(p,q)|p,q\in\bm{V}_{i}^{\prime},p\neq q\}$, where $\bm{V}=\bigcup_{i=1}^{k}\bm{V}_{i}\text{ and }\bm{V}_{i}\cap\bm{V}_{j}=\emptyset\ \forall i\neq j,\ 1\leq i,j\leq k$. Here, we use the barycenter clustering decomposition (BCD) [33] to determine the partition of $\bm{V}_{i}$. The barycenter of a route $\bm{R}=(n_{1},n_{2},\ldots,n_{L})$ is defined as $(\overline{x},\overline{y})=\frac{1}{L-1}\sum_{i=1}^{L-1}(x_{i},y_{i}),$ where $(x_{i},y_{i})$ is the Cartesian plane coordinate of node $n_{i}$. For each subproblem $\bm{G}_{i}$, the same MS is applied (lines 3-3). After all MS procedures are finished, the solutions from all subproblems are assembled to construct a complete solution $\bm{S}^{\prime}$ for the original problem.
TABLE I: Setting of algorithm parameters.
Parameter $G$ $N$ $\alpha$ $B$ $sr$ $w_{1},w_{2}$
$akb_{small}$ 20 9 3 5 $1.0$ $0.2,0.4$
$akb_{medium}$ 20 4 3 5 $0.5$ $0.1,0.2$
$jd$ 20 4 3 5 $0.1$ $0.05,0.05$
V Experiments
To evaluate the effectiveness of HMA, we compared it with the recent algorithms for solving EVRP-TW-SPD, Adapt-CMSA-STD [11] and Adapt-CMSA-SETCOV [17], using the existing benchmark set. Then, we introduced a new benchmark set derived from a real-world application that includes large-scale instances. The evaluation results of HMA on the new benchmark were also reported. Finally, we conducted a comprehensive ablation study to assess the effectiveness of PSSI, CDNS, and the hybrid search framework. The source code of HMA and the new benchmark set are open-sourced anonymously at https://anonymous.4open.science/r/EVRP-TW-SPD-HMA-code-dataset.
TABLE II: Comparative results on small-scale instances in the akb set.
CPLEX Adapt-CMSA-STD Adapt-CMSA-SETCOV HMA
Instance m best time m best avg time m best avg time m best avg time
c101C5 2 2257.75 0.43 2 2257.75 2257.75 0.02 2 2257.75 2257.75 0.01 2 2257.75 2257.75 0.38
c103C5 1 1175.37 0.41 1 1175.37 1175.37 0.54 1 1175.37 1175.37 0.68 1 1175.37 1175.37 0.21
c206C5 1 1242.56 0.58 1 1242.56 1242.56 0.03 1 1242.56 1242.56 0.00 1 1242.56 1242.56 0.19
c208C5 1 1158.48 0.08 1 1158.48 1158.48 0.01 1 1158.48 1158.48 0.00 1 1158.48 1158.48 0.24
r104C5 2 2136.69 0.02 2 2136.69 2136.69 0.07 2 2136.69 2136.69 0.01 2 2136.69 2136.69 0.93
r105C5 2 2156.08 0.03 2 2156.08 2156.08 0.01 2 2156.08 2156.08 0.00 2 2156.08 2156.08 1.28
r202C5 1 1128.78 0.06 1 1128.78 1128.78 9.00 1 1128.78 1128.78 0.00 1 1128.78 1128.78 0.24
r203C5 1 1179.06 0.03 1 1179.06 1179.06 0.22 1 1179.06 1179.06 0.05 1 1179.06 1179.06 0.17
rc105C5 2 2233.77 2.18 2 2233.77 2233.77 0.09 2 2233.77 2233.77 0.04 2 2233.77 2233.77 0.86
rc108C5 2 2253.93 0.19 2 2253.93 2253.93 0.01 2 2253.93 2253.93 0.00 2 2253.93 2253.93 0.58
rc204C5 1 1176.39 0.25 1 1176.39 1176.39 0.07 1 1176.39 1176.39 0.01 1 1176.39 1176.39 0.21
rc208C5 1 1167.98 0.12 1 1167.98 1167.98 0.52 1 1167.98 1167.98 0.03 1 1167.98 1167.98 0.19
c101C10 3 3388.25 76.74 3 3388.25 3388.25 0.28 3 3388.25 3388.55 0.33 3 3388.25 3388.25 1.70
c104C10 2 2273.93 2.16 2 2273.93 2273.93 0.48 2 2273.93 2273.93 29.02 2 2273.93 2273.93 1.10
c202C10 1 1304.06 6.34 1 1304.06 1304.06 0.45 1 1304.06 1304.06 0.10 1 1304.06 1304.06 1.40
c205C10 2 2228.28 0.08 2 2228.28 2228.28 10.82 2 2228.28 2228.28 0.03 2 2228.28 2228.28 0.45
r102C10 3 3249.19 0.56 3 3249.19 3249.19 18.22 3 3249.19 3249.19 0.01 3 3249.19 3249.19 1.03
r103C10 2 2206.12 17.49 2 2206.12 2206.12 5.27 2 2206.12 2206.12 22.01 2 2206.12 2206.12 3.38
r201C10 1 1241.51 61.20 1 1241.51 1241.51 22.71 1 1241.51 1241.51 15.97 1 1241.51 1241.51 3.29
r203C10 1 1218.21 1.57 1 1218.21 1218.21 14.06 1 1218.21 1218.21 27.17 1 1218.21 1218.21 0.97
rc102C10 4 4423.51 0.15 4 4423.51 4423.51 0.24 4 4423.51 4423.51 0.01 4 4423.51 4423.51 3.27
rc108C10 3 3345.93 3.72 3 3345.93 3345.93 14.61 3 3345.93 3345.93 0.01 3 3345.93 3345.93 1.29
rc201C10 1 1412.86 22756.49 1 1412.86 1412.86 1.97 1 1412.86 1412.86 0.93 1 1412.86 1412.86 6.03
rc205C10 2 2325.98 0.13 2 2325.98 2325.98 0.13 2 2325.98 2325.98 0.45 2 2325.98 2325.98 0.74
c103C15 3 3348.46 5046.46 3 3348.46 3348.47 48.19 3 3348.46 3348.46 3.76 3 3348.46 3348.46 3.30
c106C15 3 3275.13 0.90 3 3275.13 3275.13 4.56 3 3275.13 3275.13 32.34 3 3275.13 3275.13 2.42
c202C15 2 2383.62 43.74 2 2383.62 2383.62 12.73 2 2383.62 2393.39 5.27 2 2383.62 2383.62 1.67
c208C15 2 2300.55 3.25 2 2300.55 2300.55 1.09 2 2300.55 2300.55 8.59 2 2300.55 2300.55 1.36
r102C15 5 5412.78 5046.62 5 5412.78 5412.78 1.82 5 5412.78 5412.78 0.09 5 5412.78 5412.78 4.27
r105C15 4 4336.15 5.34 4 4336.15 4336.15 1.03 4 4336.15 4336.15 0.86 4 4336.15 4336.15 2.03
r202C15 2 2361.51 5045.36 2 2358.00 2364.90 22.71 1 1507.32 1677.46 45.16 2 2358.00 2358.00 4.00
r209C15 1 1313.24 3088.27 1 1313.24 1313.24 12.23 1 1313.24 1313.24 10.98 1 1313.24 1313.24 2.83
rc103C15 4 4397.67 245.61 4 4397.67 4397.67 0.24 4 4397.67 4397.67 0.21 4 4397.67 4397.67 2.73
rc108C15 3 3370.25 822.47 3 3370.25 3370.25 41.34 3 3370.25 3370.25 0.24 3 3370.25 3370.25 1.87
rc202C15 2 2394.39 603.76 2 2394.39 2394.39 0.48 2 2394.39 2394.39 19.65 2 2394.39 2394.39 2.72
rc204C15 1 1403.38 5046.60 1 1382.22 1385.72 47.83 1 1382.22 1382.55 50.49 1 1382.22 1384.83 19.75
avg 2.06 2338.38 1331.37 2.06 2337.70 2337.99 8.17 2.03 2314.07 2319.08 7.62 2.06 2337.70 2337.77 2.20
V-A Experimental Setup
V-A1 Benchmark
As noted by Akbay et al.[11], the existing benchmark used is based on E-VRPTW instances from Schneider et al.[12], with customer demands separated into delivery and pickup demands using the method in [34]. We refer to this benchmark as the akb (authors’ initials) set. This akb set has 92 instances, including 36 small-scale instances (twelve 5-customer instances, twelve 10-customer instances, and twelve 15-customer instances, each with 2 to 8 stations) and 56 medium-scale instances (100-customer instances with 21 stations ). These instances are categorized into three groups based on the spatial distribution of customer locations: clustered (prefixed with c), randomly distributed (prefixed with r), and a mixture of random and clustered distributions (prefixed with rc). Each group is further divided into two sub-classes: type 1 features narrow time windows and small vehicle and battery capacities, while type 2 has wide time windows and large vehicle and battery capacities. All these instances are defined in a two-dimensional Euclidean space. That is, each node $i$ has the Cartesian coordinate $(x_{i},y_{i})$. The distance between two nodes $i$ and $j$ is given by the Euclidean distance $\sqrt{(x_{i}-x_{j})^{2}+(y_{i}-y_{j})^{2}}$, which is a special case of the problem formulation provided in this paper (see Section III). All distances and time between two nodes $i$ and $j$ satisfy the triangle inequality. Following [11, 17], in the experiments conducted on the akb set, we set $\mu_{1}=1000$ and $\mu_{2}=1$.
As previously mentioned, the akb set includes only synthetic instances of small and medium scales. To address the gap in benchmark instance scales, we introduced a new benchmark set derived from JD Logistics’ distribution system. The raw data collected from this system includes requests over a period in Beijing, involving both deliveries of purchased goods and collections picked up from customers. We sampled from this data to generate instances with 200, 400, 600, 800, and 1000 customers, each with 100 stations. For each problem scale, we generated 4 instances, resulting in a set of 20 large-scale instances, referred to as the jd set. Since the jd dataset is derived from real-world data, each instance provides the longitude and latitude $(lng_{i},lat_{i})$ of each node $i$, in degrees. Additionally, the distances and time between two nodes $i$ and $j$ might not satisfy the triangle inequality, which represents a general real-world case of the problem formulation. Unlike the akb set, $\mu_{1}=300$ and $\mu_{2}=0.014$ in the jd set are directly given based on the estimated values in the application.
V-A2 Parameter Setting
We determined the parameter settings of HMA through preliminary experiments. The settings for different problem scales are detailed in Table I. Specifically, HMA requires setting six parameters: $G$, $N$, $\alpha$, $B$, $sr$, and $w_{1},w_{2}$. $G$ denotes the longest consecutive generation without improvement in HMA, and $N$ is the population size in MS (see Algorithm 3). $\alpha$ and $B$ refer to the proportionality parameter and the maximum number of generations in PSI, respectively. $sr$ specifies the selection range in station ranking list (see Algorithm 1). $w_{1}$ and $w_{2}$ are the lower and upper bounds for the proportion of customers removed in the destroy-repair operator (see Algorithm 3). All the experiments were run on an Intel Xeon Gold 6240 machine with 377 GB RAM and 32 cores (2.60 GHz, 25 MB Cache), running Rocky Linux 8.5. HMA was implemented in C++.
TABLE III: Comparative results on medium-scale instances in the akb set.
Adapt-CMSA-STD Adapt-CMSA-SETCOV HMA
Instance m best avg time m best avg time m best avg time
c101 12 13043.40 13043.42 270.56 12 13057.80 13063.54 205.53 12 13043.38 13043.43 12.57
c102 11 12056.80 12920.23 393.95 11 12073.10 12944.34 329.34 10 11089.80 12033.33 83.87
c103 11 12004.70 12026.90 317.63 10 11134.90 11917.80 504.47 10 10945.45 10945.48 136.32
c104 10 10872.80 11353.78 442.55 10 10870.70 10876.49 427.43 9 9890.98 10767.31 231.74
c105 11 12023.80 12341.60 394.88 11 12034.10 12068.86 409.38 11 12030.81 12130.20 64.75
c106 11 12013.10 12438.06 458.04 11 12025.70 12059.29 305.45 10 11040.32 11825.37 73.50
c107 11 12006.40 12023.97 378.24 11 12026.70 12046.38 276.09 10 11021.16 11815.55 67.93
c108 11 11994.70 12016.10 407.11 10 11025.80 11822.60 391.00 10 11006.10 11597.49 137.57
c109 10 11042.20 11885.30 502.22 10 10941.00 11180.77 524.19 10 10911.47 10936.41 184.63
c201 4 4629.95 4629.95 26.41 4 4678.37 4703.43 274.65 4 4629.95 4629.95 9.55
c202 4 4629.95 4629.95 192.19 4 4664.26 4706.94 277.38 4 4629.95 4629.95 34.02
c203 4 4632.27 4690.06 520.20 4 4641.45 4734.31 349.57 4 4629.95 4629.95 47.47
c204 4 4633.08 4665.78 563.25 4 4660.64 4737.07 503.66 4 4628.91 4628.91 102.73
c205 4 4629.95 4629.95 54.00 4 4629.95 4629.95 88.12 4 4629.95 4629.95 31.58
c206 4 4629.95 4629.95 149.83 4 4629.95 4629.95 142.64 4 4629.95 4630.26 39.25
c207 4 4629.95 4629.95 179.74 4 4629.95 4635.27 182.89 4 4629.95 4630.89 62.34
c208 4 4629.95 4629.95 200.06 4 4629.95 4629.95 183.86 4 4629.95 4629.95 50.89
r101 18 19633.80 19939.79 459.20 18 19640.60 19661.15 476.75 17 18636.96 19512.43 75.46
r102 17 18470.80 19292.16 496.96 16 17474.10 17696.35 561.10 16 17424.75 17526.29 115.80
r103 15 16296.50 17050.75 499.65 14 15280.30 15306.17 449.49 13 14209.21 14721.75 187.46
r104 13 14141.10 14255.53 433.34 12 13084.30 13111.31 538.23 11 12051.19 12066.53 285.75
r105 15 16389.20 17212.83 477.88 14 15471.30 16346.10 429.92 14 15366.33 16159.23 146.98
r106 15 16292.00 16836.67 492.99 14 15314.80 15441.68 524.23 13 14265.55 14373.31 222.14
r107 13 14168.90 15016.67 478.40 12 13140.10 13669.50 550.64 12 13112.66 13118.04 226.39
r108 12 13079.80 13531.30 468.60 11 12073.70 12998.17 521.61 11 12016.74 12025.10 289.00
r109 14 15237.30 15674.51 533.66 13 14220.80 14468.12 523.07 12 13203.74 13993.94 237.95
r110 13 14170.20 14905.73 370.99 12 13114.30 13544.76 541.43 11 12070.04 12482.82 270.95
r111 12 13144.20 14584.19 489.45 12 13148.80 13965.98 503.08 12 13070.64 13077.15 225.63
r112 12 13155.60 14053.56 331.29 12 13044.10 13078.65 597.35 11 12001.79 12012.62 233.83
r201 4 5192.33 5216.92 506.12 4 5276.75 5363.04 131.64 3 4268.88 4985.55 264.02
r202 3 4250.70 5020.88 483.79 3 4193.33 4940.22 615.58 3 4054.75 4063.20 363.02
r203 3 3942.74 4352.52 609.82 3 3985.02 4060.18 577.82 3 3898.68 3910.77 541.38
r204 3 3820.72 3854.31 553.01 3 3793.76 3827.81 615.67 2 2815.58 3544.13 438.26
r205 3 4055.28 4124.64 514.20 3 4065.06 4126.45 253.23 3 3995.35 4003.09 512.99
r206 3 3978.10 4065.05 531.38 3 3991.44 4047.16 550.82 3 3925.95 3938.12 600.35
r207 3 3878.91 3910.07 463.55 3 3881.97 3918.29 616.95 2 2851.75 3534.37 494.20
r208 3 3791.27 3829.39 596.95 3 3732.80 3776.20 629.30 2 2740.06 2756.80 589.42
r209 3 3975.64 4015.90 467.75 3 3933.55 3977.24 538.01 3 3877.24 3885.71 636.14
r210 3 3920.37 3984.78 530.97 3 3926.79 3961.42 520.11 3 3846.62 3854.33 555.78
r211 3 3814.42 3893.38 580.24 3 3824.47 3857.62 561.62 3 3772.67 3789.62 477.61
rc101 16 17667.70 18513.67 504.48 16 17696.20 17741.63 442.05 15 16661.53 17462.90 152.22
rc102 16 17576.80 17909.78 392.20 15 16558.20 16628.46 422.70 14 15510.31 15618.29 256.61
rc103 14 15366.90 16245.06 537.04 13 14358.20 14999.16 485.96 12 13343.56 13648.08 297.57
rc104 13 14270.50 14315.17 447.54 12 13222.50 13261.65 490.66 11 12150.47 12168.35 388.45
rc105 15 16500.90 16933.42 458.25 14 15470.70 15820.90 449.50 14 15447.52 15454.71 164.64
rc106 14 15432.50 16069.05 444.10 13 14448.30 15249.17 406.17 13 14388.12 14609.85 190.86
rc107 13 14313.40 14437.31 540.67 12 13276.50 13386.51 518.89 11 12234.76 13048.02 232.42
rc108 12 13226.00 13891.71 435.95 12 13184.70 13214.50 504.05 11 12140.28 12152.42 271.54
rc201 4 5504.77 5819.06 494.52 4 5617.75 5786.48 226.81 4 5444.90 5450.77 254.36
rc202 4 5324.64 5442.41 416.84 4 5436.63 5541.80 138.34 4 5240.40 5249.87 192.66
rc203 4 5109.88 5177.69 452.57 4 5086.44 5159.15 531.94 3 4078.55 4100.32 597.82
rc204 3 4036.49 4525.03 523.68 3 3962.88 4252.28 631.65 3 3892.97 3902.06 528.96
rc205 4 5260.14 5338.50 426.74 4 5285.41 5375.54 221.24 4 5149.13 5152.79 368.18
rc206 4 5234.55 5289.90 470.82 4 5210.57 5275.72 395.42 3 4200.69 4323.33 517.29
rc207 3 4150.60 4930.81 488.10 3 4197.81 4650.01 607.24 3 3979.65 4019.57 490.17
rc208 3 3977.50 4046.09 535.72 3 3920.17 4002.79 527.17 3 3838.78 3857.49 457.65
avg 8.48 9568.86 9905.20 435.54 8.18 9266.06 9479.93 432.20 7.77 8807.09 9030.14 262.90
V-B Results on Small-scale and Medium-scale Instances
We compared HMA with two recently proposed algorithms, Adapt-CMSA-STD and Adapt-CMSA-SETCOV, on the akb set. Since their source code is not publicly available, we obtained the testing results directly from the original publications [11, 17]. To ensure a fair comparison, we rescaled the computation time to a common measure considering the CPUs used. Specifically, we rescaled the times based on the PassMark score of the CPUs used in the studies relative to our 2.60 GHz Intel Xeon Gold 6240. The PassMark score, as the performance of a single core of the respective CPU, for the 2.93 GHz X5670 used in [11, 17] is 1398, while the score of our CPU is 1990. The rescaled time in seconds is given in Tables II and III. Additionally, the run time limit on the small-scale and medium-scale instances is rescaled to 105 and 630 seconds.
Tables II and III are structured as follows. The first column lists the instance names. The “m” columns indicate the number of vehicles used in the respective solutions. For Adapt-CMSA-STD, Adapt-CMSA-SETCOV, and HMA, the “best” column refers to the $TC$ of the best solution found in 10 independent runs, while the “avg’ column refers to the average $TC$ of the 10 runs. Finally, the “time” column shows the computation time (in seconds) for CPLEX and the average computation time for Adapt-CMSA-STD, Adapt-CMSA-SETCOV, and HMA to find the best solutions in each run, respectively. For each instance, the best performance w.r.t. “best” and “avg” are highlighted in gray, respectively.
TABLE IV: Comparative results on large-scale instances in the jd set with a short time limit.
HMA w/o BCD HMA
Instance m best avg±std time m best avg±std time p-value
jd200_1 44 71043.42 72243.96±614.67 1884.42 44 70979.75 72348.31±724.19 1821.75 0.6501
jd200_2 39 59660.84 60785.77±551.51 1892.70 39 59460.27 60688.60±697.48 1825.64 0.7624
jd200_3 44 66634.23 67043.56±269.90 1872.56 43 66254.27 66999.61±526.33 1839.70 0.8798
jd200_4 39 60988.30 62033.06±488.03 1873.78 39 60941.60 61753.23±523.17 1843.03 0.2568
jd400_1 79 116029.37 117199.07±954.98 5784.07 79 115981.66 116987.12±952.18 5443.33 0.4963
jd400_2 77 116527.33 117584.91±736.78 5734.43 77 116205.05 117098.09±563.09 5430.44 0.1041
jd400_3 76 109030.69 109934.10±508.53 5576.75 76 108441.53 109140.65±451.85 5421.66 0.0032
jd400_4 83 128988.11 129937.62±710.96 5761.35 83 128894.04 129654.11±858.13 5428.58 0.1509
jd600_1 112 167369.71 169339.19±1098.95 9351.29 112 167297.70 168836.18±1087.42 9033.06 0.3258
jd600_2 113 171996.76 174044.48±1610.00 9643.47 113 171512.97 173480.89±1384.65 9019.82 0.3447
jd600_3 112 154336.98 155381.50±773.48 9610.60 111 153092.64 154325.16±840.34 9037.23 0.0191
jd600_4 104 150513.45 151553.99±649.75 9753.66 103 150123.89 150975.72±609.08 9049.35 0.0343
jd800_1 147 216364.67 217948.77±1027.41 13710.86 146 214710.04 216810.78±1542.74 12625.32 0.0757
jd800_2 148 226218.76 228162.86±1987.21 13523.34 149 223610.37 226527.29±2364.16 12619.90 0.0963
jd800_3 148 206772.88 208548.00±1067.70 13401.35 147 205215.92 207375.35±1432.89 12619.30 0.0494
jd800_4 136 196076.58 197143.07±637.84 13603.29 134 194004.95 196257.28±1325.29 12621.05 0.0696
jd1000_1 179 261338.90 262810.90±1082.93 17483.69 178 257410.81 260877.67±2210.64 16224.57 0.0343
jd1000_2 183 276597.69 279318.78±1341.59 16643.43 184 274079.42 276515.30±1881.23 16224.80 0.0028
jd1000_3 184 256229.79 258040.35±1467.13 16631.34 182 252772.63 256102.57±2369.95 16221.87 0.0494
jd1000_4 169 243354.91 244808.00±1235.27 17552.66 168 240559.58 242952.46±2044.65 16218.06 0.0284
Based on the numerical results in Table II, the following observations can be made. On small-scale instances, CPLEX, Adapt-CMSA-STD, Adapt-CMSA-SETCOV, and HMA are comparable, showing no significant difference in their best or average values due to the small solution space. Moreover, it is worth mentioning that HMA exhibits minimal computation time variation as the problem scale increases, with an average computation time of 2.20 seconds, which is much shorter than that of the other three algorithms.
Generally, the performance of algorithms on instances of larger scales is of more interest. Overall, HMA is the best-performing algorithm on medium-scale instances as shown in Table III. In the $\sim$ and $\sim$ subsets, the best and average solution quality obtained by HMA are consistently better than those of Adapt-CMSA-STD and Adapt-CMSA-SETCOV. Note that for the akb set, minimizing the number of the used EVs has a higher priority than minimizing $TD$. It is worth mentioning that HMA obtained competitive solutions with fewer EVs in 23 out of 56 instances. More specifically, these instances include 4 out of 9 in the c1$\sim$ subset, 7 out of 12 in the r1$\sim$ subset, 4 out of 11 in the r2$\sim$ subset, 6 out of 8 in the rc1$\sim$ subset, and 2 out of 8 in the rc2$\sim$ subset. Such results indicate that HMA performed both strongly and robustly. Moreover, HMA required significantly less computation time, taking only about 60% of the time compared to the other two algorithms. In summary, the observations demonstrate that HMA outperforms all recent EVRP-TW-SPD algorithms across a broad range of problem instances, and in particular, it has significantly raised the performance bar on the akb set.
TABLE V: Comparative results on large-scale instances in the jd set with a long time limit.
HMA w/o BCD HMA
Instance m best avg±std time m best avg±std time p-value
jd200_1 44 70740.90 71335.98±403.87 3678.82 44 70440.57 71409.68±512.27 3510.36 0.6501
jd200_2 39 59389.70 60004.86±419.68 3661.48 38 58530.22 59823.10±570.84 3612.10 0.3643
jd200_3 42 65201.82 65844.88±445.08 3544.28 43 65812.71 66206.45±284.52 3534.71 0.0588
jd200_4 39 60747.28 60983.75±250.76 3571.75 39 60815.01 61008.17±184.95 3362.58 0.4497
jd400_1 78 114537.34 116518.79±1422.92 10492.89 78 114489.28 115579.17±715.05 10474.83 0.1306
jd400_2 76 115303.21 116803.18±612.69 10409.36 77 114668.85 115658.43±606.08 10815.96 0.0015
jd400_3 76 106951.59 107842.05±588.35 10547.36 76 107600.70 107927.90±351.95 10063.53 0.8206
jd400_4 81 126613.16 127573.34±914.21 10687.95 82 126927.71 127814.74±492.32 10362.35 0.2265
jd600_1 112 166524.71 168729.52±1110.96 14626.36 112 166113.91 166952.86±681.25 18016.97 0.0019
jd600_2 113 171996.76 173256.93±1387.16 14450.07 113 170362.24 171446.42±618.81 17416.13 0.0007
jd600_3 110 151879.40 153745.61±1118.40 16444.28 109 151185.33 152554.96±942.57 17806.93 0.0494
jd600_4 103 149893.97 151118.39±658.62 13495.48 103 148729.43 149447.53±563.44 17432.51 0.0004
jd800_1 147 216194.15 217442.46±805.04 20030.45 146 213065.47 214228.94±974.71 25217.34 0.0002
jd800_2 148 226218.76 227630.94±1710.11 19306.58 147 222333.51 223685.24±1338.04 25214.00 0.0007
jd800_3 147 205429.64 207285.21±1202.21 23965.29 146 203751.04 205483.36±1355.49 24410.31 0.0082
jd800_4 135 194738.56 196802.01±952.89 23059.46 134 192848.13 194144.22±987.70 24949.64 0.0005
jd1000_1 178 260502.17 262307.88±1199.71 29046.07 178 256626.16 257932.03±1114.15 32159.10 0.0002
jd1000_2 182 275972.37 278824.43±1564.13 28589.01 181 270788.78 273195.12±1537.57 32421.83 0.0002
jd1000_3 184 256229.79 257553.76±1262.58 25397.05 181 250802.03 253040.02±1411.33 31452.40 0.0002
jd1000_4 169 242662.53 244305.00±1250.99 27539.25 167 238562.78 239966.65±1244.47 32276.48 0.0002
V-C Results on Large-scale Instances
To demonstrate the effectiveness of HMA for solving large-scale instances in the $jd$ set, we limited both short and long maximum computation time and compared HMA with HMA without the BCD decomposition (denoted HMA w/o BCD). Here, we omit the existing algorithms [11, 17] for solving EVRP-TW-SPD because they are not open-sourced. Specifically, instances with 200, 400, 600, 800, and 1000 customers were decomposed into 2, 4, 6, 8, and 10 subproblems in the MS of HMA (see Algorithm 3). The short time limits were set to 1800, 5400, 9000, 12600, and 16200 seconds (see Table IV), while the long time limits were 3600, 10800, 18000, 25200, and 32400 seconds (see Table V). Note that BCD requires the Cartesian coordinate of node $n_{i}$, while the $jd$ set contains longitude and latitude. Since the area covered is not large, we can use Mercator Projection to convert $(lng_{i},lat_{i})$ into $(x_{i},y_{i})$, specifically, $x_{i}=R\cdot\frac{\pi\cdot lng_{i}}{180}$, $y_{i}=R\cdot\ln\left(\tan\left(\frac{\pi}{4}+\frac{\pi\cdot lat_{i}}{360}\right)\right)$, where $R$ is the semi-major axis of the WGS84 ellipsoid model. In Tables IV and V, the “avg±std” columns represent the average and standard deviation of $TC$ for the 10 best solutions found in 10 independent runs. We used the Wilcoxon rank-sum test with significance level $p=0.05$ to compare the performance of HMA w/o BCD and HMA, with p-values shown in the last column. The best performance w.r.t. “best” is highlighted in gray, and w.r.t. “avg±std” is highlighted in bold if it is significantly better than the other algorithm, respectively.
From Table IV, it can be observed that under a short time limit, HMA consistently found better solutions on all instances of the $jd$ set compared to HMA w/o BCD. The average performance of HMA is significantly better as the problem scale increases. In Table V, under a long time limit, HMA w/o BCD performed reversely better on some instances with 200 and 400 customers. This is due to the challenge of designing an ideal decomposition, which would partition all customers belonging to the same route in the optimal solution into the same subproblem. However, other decompositions, such as BCD, solve subproblems efficiently but may result in suboptimal solutions. Nonetheless, as the problem scale grows, the decomposition strategy still shows an obvious advantage on instances of larger scale with 600, 800, and 1000 customers since HMA obtained significantly higher-quality solutions. The performance gap between HMA w/o BCD and HMA across 10 independent runs, measured as $[TC_{1}-TC_{2}]/{TC}_{2}$, where $TC_{1}$ and $TC_{2}$ are the value of best-found solutions from HMA w/o BCD and HMA, is 0.61%, 1.26%, and 1.83% averaged on instances with 600, 800, and 1000 customers, respectively. As the problem scale grows, the performance gap between HMA w/o BCD and HMA also becomes larger.
Figure 2: Comparative effects of PSI and SSI on each subset. The outer ring represents the exclusive contribution ratio (ECR) of SSI and PSI, with the remainder being shared contributions; The inner ring shows the time allocation ratio (TAR) of SSI and PSI, with the remainder representing others’ TAR.
V-D Effectiveness of PSSI, CDNS, and Hybrid Framework
First, we separately analyzed the comparative effects of PSI and SSI for the following reasons: PSSI selects the best result from PSI and SSI. Therefore, we are more concerned about which strategy, PSI or SSI, achieves better station insertion on a route. Note that SSI produces only a single route, compared to PSI with a population. We need to validate whether PSI can really find other higher-quality station insertion configurations with acceptable additional computational cost. To this end, we ran HMA once on all instances, keeping track of the performance of PSI and SSI in terms of station insertion on the same route. Figure 2 presents the performance of PSI and SSI: the exclusive contribution ratio (ECR) is computed as $n_{1}/n_{2}$ for each instance, where $n_{1}$ is the number of insertion results obtained exclusively from SSI or PSI and $n_{2}$ is the total number of insertion results; the time allocation ratio (TAR) is computed as $t_{1}/t_{2}$ for each instance, where $t_{1}$ is the cumulative time consumed in SSI or PSI components, $t_{2}$ is the total time consumed in HMA. Note that the outer ring “ECR” and inner ring “TAR” in Figure 2 are the average values on each subset.
From the results, it can be clearly seen that PSI can really find other higher-quality station insertions, as ECR $>0$ indicates that it found some strictly better insertions than SSI. In the $jd$ set, the ECR of PSI even achieved 46.90%, accounting for a large proportion of contributions to PSSI. Notably, although PSI with a genetic algorithm took approximately $2\sim 8$ times runtime than SSI, this additional runtime is worthwhile and acceptable, as SSI cannot further improve its insertion results. Thus, combining the results of both PSI and SSI can better avoid local optima and construct higher-quality routes with acceptable additional computational cost.
Second, we conducted an ablation study to evaluate the effectiveness of other components in HMA. We tested the following four variants on the $akb$ and $jd$ sets, each of which is different from HMA in one component:
  1. 1.
    w/o CLS: lines 2-2 of Algorithm 2 are removed;
  2. 2.
    w/o ALS: lines 2-2 of Algorithm 2 are removed;
  3. 3.
    w/o MS: lines 3-3 of Algorithm 3 are removed;
  4. 4.
    w/o LNS: lines 3-3, 3-3 of Algorithm 3 are removed.
The first two variants were to validate the effectiveness of CDNS, while the latter two were to validate the effectiveness of the hybrid framework. For each variant, the performance degradation ratio (PDR) is computed as $[TC_{1}-TC_{2}]/{TC}_{2}$ for each instance, where $TC_{1}$ and $TC_{2}$ are the average $TC$ obtained by the variant and HMA across 10 independent runs, respectively. PDR $>0$ indicates the component is useful for this specific instance. Larger values indicate a greater impact. Table VI presents the average PDR on each subset.
TABLE VI: Average PDR for the four HMA variants on each subset.
Subset w/o CLS w/o ALS w/o MS w/o LNS
$akb_{small}$ 1.80% 0.00% 2.94% 0.00%
$akb_{medium}$ 1.03% 0.06% 3.49% 0.37%
$jd$ 0.00% 4.03% 0.60% 2.47%
Avg.PDR 0.84% 0.77% 2.98% 0.75%
The results show that, on average, no component negatively affects the performance of HMA. Specifically, the “w/o CLS” and “w/o ALS” results show that CLS benefits small-scale and medium-scale instances, while ALS is more effective for large-scale instances. CLS is good at refined search, which is beneficial for small solution spaces. In contrast, ALS leverages existing efficient VRP-TW-SPD move evaluation, advantageous for large solution spaces that require fast exploration. Similarly, the “w/o MS” and “w/o LNS” results indicate that MS is more effective for small-scale and medium-scale instances due to population diversity, whereas a fast search of LNS benefits large-scale instances. The population maintains solution diversity in smaller problems; however, as the problem scale increases, the dominant role of individual-based search capabilities becomes more evident.
VI Conclusion
This work investigated an emerging routing problem, namely EVRP-TW-SPD, and proposed a hybrid memetic algorithm called HMA to solve it. HMA integrates two novel components: PSSI for handling partial recharges and CDNS as the local search procedure. Both components can be easily applied to other EVRP variants. Comprehensive experiments showed that HMA outperformed existing methods across a wide range of instances. Moreover, a new, large-scale benchmark set was established from a real-world JD logistics application. This set contains 20 problem instances with 200, 400, 600, 800, and 1000 customers, each with 100 stations,
HMA can be further improved in several directions. First, the PSSI procedure could benefit from a learned function (e.g., as a neural network [35]) that evaluates insertions instead of heuristic rules, which might lead to higher-quality feasible routes within even smaller time budget. Second, dividing the original problem into smaller, manageable subproblems enhances efficiency and yields high-quality solutions. However, the success of this approach heavily depends on the effectiveness of the decomposition strategy. Rather than relying on hand-designed strategies, future improvements could leverage learning-based methods [36] or general-purpose pretrained models [37] to automatically design effective decomposition strategies tailored to specific customer distributions.
Appendix A Proof of Proposition 1
Proof. Consider a feasible EVRP-TW-SPD solution $\bm{S}$. By definition, $\bm{S}$ satisfies all electricity, capacity, and time window constraints. Let $\bar{\bm{S}}$ be the solution obtained by removing all charging stations in $\bm{S}$. Suppose the travel time between any two nodes satisfies the triangle inequality, then for any two successive customers $c_{i}$ and $c_{j}$ in $\bar{\bm{S}}$ that are separated by a charging station $f$, the triangle inequality ensures:
$t_{c_{i}c_{j}}\leq t_{c_{i}f}+t_{fc_{j}}+g\cdot q_{f},$
where $g\cdot q_{f}$ is the charging time at station $f$. This inequality implies that the travel time between customers in $\bar{\bm{S}}$ is no greater than in $\bm{S}$. Consequently, $\bar{\bm{S}}$ satisfies all capacity and time window constraints of VRP-TW-SPD, making it a feasible solution.
In cases where the travel time does not satisfy the triangle inequality, we can preprocess the problem instance to ensure this property holds. Specifically, we first apply the Floyd-Warshall algorithm to compute the shortest path in terms of travel time between every pair of nodes. These shortest paths are allowed to pass through charging stations but not through the depot or other customers. We then reconstruct the edge set $\bm{E}$ using these shortest paths, defining each new edge (or “hyperarc”) as $(i,j)$ to be the shortest path from node $i$ to $j$. The travel time associated with these new edges naturally satisfies the triangle inequality, allowing Proposition 1 to hold for any problem instance. ∎